A random sample of n = 15 heat pumps of a certain type yielded the following observations on lifetime (in years):

2.0     1.5     6.0     1.7 5.3 0.4     1.0     5.3
15.8 0.5 4.8 0.9     12.1     5.3 0.6

(a) Assume that the lifetime distribution is exponential and use an argument parallel to that of this example to obtain a 95% CI for expected (true average) lifetime. (Round your answers to two decimal places.) ( , ) years

c) What is a 95% CI for the standard deviation of the lifetime distribution? [Hint: What is the standard deviation of an exponential random variable?] (Round your answers to two decimal places.) ( , ) years
Answer #1

solutions & Given n=15, 2= 4:21,5= 4.5103 df = n-1 = 15-1 = 14. At 95% confidence tap. = 2.145 The 95% confiducenternal for M